OpenStudy (anonymous):

The surface areas of two similar solids are 289 ft2 and 900 ft2. The volume of the larger solid is 1,755 ft3. What is the volume of the smaller solid?

OpenStudy (anonymous):

The surface areas of two similar objects are proportional to the square of their sides. Small side^2 / large side^2 = 900/289 Small side/large side = 30/17 Volume of two similar objects is proportional to the cubes of their sides. Smaller volume/1755 = 17^3/30^3 Smaller volume = 1755(4913/27000) = 319

OpenStudy (anonymous):

dang your smart!! lol. thanks again xD I'm taking notes about every work you gave me

OpenStudy (anonymous):

You're welcome!! Xo!